jhonyy9 (jhonyy9):

- prove that all even numbers from the set of natural numbers (6 till 20)can be writing like sum of p and k ,if we know that p and k are prims from the set of prim numbers (2 till 19).

OpenStudy (mr.math):

6=3+3 8=5+3 10=5+5 12=5+7 14=7+7 16=11+5 18=11+7 20=17+3

jhonyy9 (jhonyy9):

- so how you can prove this till 1000 ?

OpenStudy (mr.math):

This is a famous conjecture that has not yet been proved. Read here http://en.wikipedia.org/wiki/Goldbach's_conjecture

jhonyy9 (jhonyy9):

- so we know that all even numbers can be writing in the form of 2n when n is one natural number from the set of natural numbers N - p and k are prims so p=2n+1 and k=2n+1 - so p+k=2a+1+2b+1=2(a+b+1) so if (a+b+1)=n such that for every n there are one a and one b such that a+b+1=n so than for every 2n=m will exist one p and one k such that always p+k=m - is this true ??? ... so can being this true ???

OpenStudy (mr.math):

I wish it was true, you would win $1,000,000 and send me half of it! ;)

jhonyy9 (jhonyy9):

- so if you wish that is true and this will wish it again to much persons from All over the world - so than i think that this prove will be accepted REALLY ... - thank you very much